Operation Summary for November 7, 2011 - November 9, 2011
Monday November 7
During the day shift, at 8:31 AM, a Booster RF station (BRF15) tripped off; it reset. At 8:59 AM, BRF14 tripped off for the first of two times in 13 minutes; it reset. Kicker personnel reported at 9:05 AM that they found and started repairs on a high voltage cable for the MI abort gap kicker (KPS4C). At 9:11 AM, BRF14 and BRF15 tripped off; they reset. At 9:15 AM, BRF4, BRF14, and BRF15 tripped off for the first of four times in 10 minutes; they reset. A Booster person began repairs on a kicker (MKS09) at 9:51 AM. Operations injected protons into the Recycler around 10:11 AM. At 11:14 AM, a Linac RF station (LRF4) tripped off; Linac personnel investigated and began repairs. Booster RF personnel began repairs on BRF13. Operators tried to search and secure the Muon rings (formerly the Antiproton Source) several times, but a problem with an interlock light link stopped them; the Interlock technicians could not find a problem, so the light link interlock was bypassed and guards were posted while operators completed the search and secure. FESS personnel reported cleaning the pond water pump strainers at Tevatron sectors F2 and F3 at 12:49 PM. A NuMI BPM (NBPM65) stopped polling; operators rebooted the system at 1:13 PM. At 1:15 PM, the NuMI beam permit tripped off due to a BPM; operators rebooted the BPM system again. At 1:25 PM, LRF4 tripped off for the first of three times in 14 minutes; it reset.
During the evening shift, at 10:32 PM, the MINERvA Run Coordinator reported that the MINOS near detector had failed; Operations halted beam to NuMI. Operators loaded an energy conservation timeline into the Time Line Generator (TLG).
Tuesday November 8
The midnight shift began with NuMI off due to a problem with the MINOS near detector and with Operations monitoring beam to the MiniBooNE experiment.
During the midnight shift, LRF4 tripped off at 3:04 AM, it reset. At 3:19 AM, operators reported that the MI-30 BPM VME had stopped polling twice in two hours; operators locally rebooted the system twice. At 3:34 AM, three Linac quadrupole power supplies (L:QPS108, L:QPS13, and L:QPS15) tripped off; they reset. Controls personnel reported at 7:18 AM that the CAMAC link at AP-10 was down; they investigated and replaced a fuse for the repeater chassis. At 7:30 AM, Neutron Therapy Facility personnel began preparing to treat one patient later in the morning.
During the day shift, Operations restored beam to NuMI at 8 AM. Operators tuned the 750 KeV line at 2:01 PM. Operators tuned the Booster injection at 3:50 PM. Operators reported a short downtime due to a Linac Klystron RF station (KRF7) trip on safety relay status.
During the evening shift, due to the MINOS near detector failure, the MINERvA Run Coordinator asked Operations to halt beam to NuMI at 6:32 PM; operators loaded an energy conservation timeline into the TLG. At 8:52 PM, MINERvA requested beam to NuMI be reestablished.
Wednesday November 9
The midnight shift began with Operations sending beam to the NuMI and MiniBooNE experiments.
During the midnight shift, operators reported at 5:29 AM that two Linac quadrupole power supplies (L:Q41 and L:Q43) were suffering from large current fluctuations; operators power cycled the Q4BULK power supply. Operators tuned MI at 7:55 AM to improve beam efficiencies.
Operations Coordinator (OpsCo) Plans for Today
The plans for today includes a two-hour access, beginning at noon, for MI to fill a kicker with Fluorinert.
Comments and Suggestions
|last modified 11/09/2011 email Fermilab|